For the following reaction, Kc = 255 at 1000 K. CO (g) + Cl2 (g) ⇌ COCl2 (g) A reaction mixture initially contains a CO concentration of 0.1550 M and a Cl2 concentration of 0.179 M at 1000 K. You may want to reference (Pages 628 - 635) section 16.8 while completing this problem. What is the equilibrium concentration of CO at 1000 K? Express your answer in molarity to three significant figures.
Answer #1

x= 0.14063 Because a cannot be more than 0.1550 so, equilibirlum concentration of co=(0.1550-2) M =(0.1550-0:14063) M -0.0143